# pfl-research: simulation framework for accelerating research in Private Federated Learning

### v1

@@ 4.1 Comparing performance to alternative FL simulators
FedML (PyTorch) he2020fedml 1 1 $90.95_{\pm 0.47}$ $71.23\%_{\pm 0.00}$ 21.6$\times$
TensorFlow Federated tensorflowfederated 1 1 $113.52_{\pm 1.26}$ $70.02\%_{\pm 0.36}$ 27$\times$
TensorFlow Federated tensorflowfederated 1 10 $82.23_{\pm 0.49}$ $70.12\%_{\pm 0.35}$ 19.6$\times$
Flower (PyTorch) beutel2020flower 1 1 $86.88_{\pm 1.42}$ $68.30\%_{\pm 0.12}$ 16.4$\times$
Flower (PyTorch) beutel2020flower 1 10 $29.26_{\pm 0.90}$ $67.66\%_{\pm 0.28}$ 7$\times$
FedScale lai2022fedscale 1 1 $425.2_{\pm 17.66}$ $70.41\%_{\pm 0.00}$ 101$\times$
FedScale lai2022fedscale 1 10 $301.7_{\pm 19.22}$ $70.90\%_{\pm 0.43}$ 71.8$\times$
FLUTE garcia2022flute 1 1 $67.86_{\pm 3.36}$ $62.26\%_{\pm 0.07}$ 16.1$\times$
@@
As a second comparison, we use FLAIR, which instead emphasizes the efficiency of training both a larger model and dataset in distributed simulations. Due to development time constraints, we implemented the FLAIR setup only for the two frameworks most frequently used for publications: TFF and Flower.
Table 2: Comparison of simulating FLAIR dataset with different FL frameworks. The metric is the same as FL-F-17 C-AP in song2022flair .
@@
pfl-research (PyTorch) 2 2.16 $61.8_{\pm 0.001}$ 1$\times$
TensorFlow Federated tensorflowfederated 5 18 $62.0_{\pm 0.3}$ 7.8$\times$
Flower beutel2020flower 5 28.8 31.5  [^3] Parity in performance with TFF and pfl-research has not yet been achieved for this setup at the time of this pre-print. 13.3$\times$
Table 2 shows that pfl-research outperforms the alternatives prominently in terms of speed on the larger scale FLAIR benchmark. Relative to pfl-research, TFF is more efficient on FLAIR than on CIFAR10, and vice versa for Flower. The number of processes per GPU, $p$, was tuned to minimize wall-clock time, and the results show that pfl-research needs many fewer training processes per GPU to saturate GPU utilization. Tables 1 and 2 demonstrate that before the release of pfl-research, the comparison of framework speed greatly depended on the choice of benchmarks.
The superior performance of pfl-research can be attributed to several design choices, some of which are unique to it or not commonly employed by other FL simulators: (1) there is no serialization of models; (2) only one model is initialized and preserved on the GPU at all times; (3) model parameters are updated in-place; (4) pfl-researchdoes not simulate the topology of FL, i.e. there is no main coordinating process to gather and broadcast model updates; (5) ML framework tensors are used on the GPU end-to-end in the outer training loop; and (6) pfl-researchperforms load balancing of training users, which is important particularly on FLAIR.
@@ 4.2 Scaling distributed training
Figure 2: Speedup from scaling up distributed simulations. Left panel: Sweep number of processes per GPU on CIFAR10, StackOverflow and FLAIR benchmarks and keep the hardware resources pinned. Right panel: Sweep the number of GPUs to train the StackOverflow benchmark, repeat for 1, 3 and 5 processes per GPU. Blue lines (tracked on left y-axis) show wall-clock time. Green lines (tracked in right y-axis) show total GPU hours for the same experiments.
Figure 2 shows the wall-clock time when scaling up with distributed simulations. In pfl-research, the number of worker processes is decoupled from the number of GPUs available, allowing processes to share a GPU. Using more than one process per GPU, $p>1$, always has a positive effect, but the optimal number depends on the use-case (model size, cohort size). For CIFAR10 and StackOverflow stackoverflow2023 , we can set $p=5$ and the wall-clock time is less than $0.43$ of that with $p=1$ with same resources, while for FLAIR we can only fit a maximum of $p=3$ on the GPU and the relative speedup is smaller.
The right panel of Figure 2 shows the wall-clock time when scaling up the number of GPUs on the StackOverflow benchmark. The wall-clock time is reduced to 3 hours with just one GPU when scaling up number of worker processes to five, and the wall-clock time is reduced to 15 minutes when scaling up to 32 GPUs. Thus, pfl-research scales well with number of GPUs. Since the cohort size is fixed (400 in the StackOverflow benchmark), and users take a varying amount of time to train with a fixed number of local epochs, the rate of improvement is expected to be reduced as we increase the total number of processes because there are fewer users per process to load balance. This is revealed by the total number of GPU hours to finish (green lines), which indicate that scaling up to more than eight GPUs gets materially less efficient per GPU. Another insight from Figure 2 is that when a setup already uses a large number of GPUs, scaling up the number of processes per GPU can have an adverse effect, e.g., the cross and triangle lines in the right panel and blue line in the left panel show that wall-clock time starts increasing for more processes when we already use 32 GPUs for StackOverflow. With this in mind, it is advisable to first scale the number of processes before you start adding more GPU resources.
@@ 4.3 Benchmarks for research
Table 3: Performance of FL algorithms on the pfl-research benchmark suite without DP. All experiments are run for five times with different random seed and metrics are averaged. The abbreviations are as following: C10 stands for accuracy on CIFAR10, SO stands for perplexity on StackOverflow, FLR stands for mean average precision (mAP) on FLAIR. The last three columns are the perplexity scores for three LLM benchmark datasets: Stanford Alpaca (SA), Aya, and OpenAssistant (OA). If not mentioned, the natural or artificial non-IID partitioning is used for the dataset, and -IID means partitioning with fixed number of samples per client drawn IID. Algorithms we benchmark are FedAvg mcmahan2017communication , FedProx li2020federated , FedProx with adaptive $\mu$ denoted as AdaFedProx (Appendix C.3.3 in li2020federated ), and SCAFFOLD karimireddy2020scaffold .
@@
FedAvg 70.37 69.80 60.87 65.08 61.83 3.61 4.02 6.39
FedProx 70.69 69.79 60.88 64.45 62.02 3.61 4.02 6.39
AdaFedProx 70.39 69.61 61.01 64.53 61.99 3.61 4.02 6.40
SCAFFOLD 60.17 62.43 60.82 3.73 4.77 6.90
Table 4: Performance of FL algorithms on the pfl-research benchmark suite with central DP. The same rules and notation as in Table 3 are used. G stands for Gaussian mechanism with accounting using privacy loss distribution meiser2018tight ; doroshenko2022connect and BMF stands for banded matrix factorization mechanism choquette2024amplified .
@@
FedAvg G 68.77 66.70 69.11 65.24 61.49 3.62 4.19 6.52
FedAvg BMF 69.21 67.07 62.47 65.55 61.64 3.62 4.20 6.53
FedProx G 68.72 66.75 69.12 65.24 61.55 3.62 4.20 6.52
AdaFedProx G 68.78 66.80 69.14 65.18 61.58 3.62 4.20 6.52
SCAFFOLD G 61.31 77.71 51.64 3.79 4.76 7.00
The initial set of algorithms we benchmark is shown in Table 3 without DP and in Table 4 with central DP. The breadth of algorithms is limited at this moment, but we believe we have provided a flexible foundation for future work to expand the results in terms of algorithms, complementary datasets, better hyper-parameter tuning of algorithms for more fair comparison, and other evaluation metrics, e.g. amount of communicated bits.
@@ 5 Future work
pfl-research currently provides a small collection of concrete algorithms. Our primary goal has been to release a fast, modular and easy-to-use simulator which is useful for advancing research in this area. We will continuously benchmark new algorithms and methods and provide implementations for the ones that consistently beat our baselines, and will rely heavily on the community for contributing implementations of new research.
Integrating datasets is also an ongoing effort. We plan to carefully curate more benchmarks while maintaining a diverse set of domains for the benchmarks. All benchmark results reported in Section 4.3 used PyTorch. We plan to release similar benchmarks in TensorFlow. We recognize that all our current benchmarks use cross-device FL. Cross-silo FL, where a small number of entities each maintain a larger dataset and want to collaborate in training with a few other entities, can be simulated with pfl-research, but we do not provide benchmarks for this scenario yet.  [^4] We refer to https://github.com/apple/pfl-research/blob/main/pfl/data/sampling.py#L143 for more information about cross-silo simulation in pfl-research.
We currently support only synchronous FL simulations using replicated processes. The aggregator interface can be extended to make a client-server architecture with a dedicated aggregator and coordinator process, which might be required for testing certain research ideas. We also do not provide interfaces for implementing PFL with real edge devices. This is outside the scope of pfl-research at this time.
@@ Appendix A Preliminaries
Appendix B System design
B.1 Data flow
Algorithm 1 Generalized federated learning.
@@ B.1 Data flow
Algorithm 1 describes the generalized federated learning procedure implemented in pfl-research that can run a concrete FL algorithm with a generic model and orthogonal features like aggregation and DP mechanisms.
An algorithm run for a number of central iterations, which is either determined by the algorithm or interrupted by a callback. A context $c_{i}\in\mathcal{C}$ is a recipe for how to run the next central iteration. It contains parameters for local model training and evaluation, state of the algorithm, dataset target criteria and parameters for other steps of the algorithm. A federated algorithm $alg$ operates on a group of federated datasets $\mathcal{F}$ and each context $c_{i}$ targets one federated dataset $D\in\mathcal{F}$.

### v2

@@ 4.1 Comparing performance to alternative FL simulators
FedML (PyTorch) [34] 1 1 $90.95_{\pm 0.47}$ $71.23\%_{\pm 0.00}$ 21.6$\times$
TensorFlow Federated [30] 1 1 $113.52_{\pm 1.26}$ $70.02\%_{\pm 0.36}$ 27$\times$
TensorFlow Federated [30] 1 10 $82.23_{\pm 0.49}$ $70.12\%_{\pm 0.35}$ 19.6$\times$
Flower (PyTorch) [11] 1 1 $86.88_{\pm 1.42}$ $68.30\%_{\pm 0.12}$ 16.4$\times$
Flower (PyTorch) [11] 1 10 $29.26_{\pm 0.90}$ $67.66\%_{\pm 0.28}$ 7$\times$
FedScale [50] 1 1 $425.2_{\pm 17.66}$ $70.41\%_{\pm 0.00}$ 101$\times$
FedScale [50] 1 10 $301.7_{\pm 19.22}$ $70.90\%_{\pm 0.43}$ 71.8$\times$
FLUTE [28] 1 1 $67.86_{\pm 3.36}$ $62.26\%_{\pm 0.07}$ 16.1$\times$
@@
As a second comparison, we use FLAIR, which instead emphasizes the efficiency of training both a larger model and dataset in distributed simulations. Due to development time constraints, we implemented the FLAIR setup for the two frameworks that are most frequently used for publications: TFF and Flower.
Table 2: Comparison of simulating FLAIR dataset with different FL frameworks. The metric is the same as FL-F-17 C-AP in [79]. Note that version 0.2.0 was released after the initial open-source date of pfl-research, which implements improved scheduling for training users, see Appendix B.6.
@@
pfl-research 0.1.0 (PyTorch) 2 2.16 $61.8_{\pm 0.001}$
pfl-research 0.2.0 (PyTorch) 2 1.77 $61.1_{\pm 0.3}$
pfl-research 0.2.0 (PyTorch) with central DP 2 1.93 - -
TensorFlow Federated [30] 5 18 $62.0_{\pm 0.3}$ 10.2$\times$
Flower [11] 5 30.5 $62.2_{\pm 0.1}$ 17.2$\times$
Table 2 shows that pfl-research outperforms the alternatives prominently in terms of speed on the larger scale FLAIR benchmark. Relative to pfl-research, TFF is more efficient on FLAIR than on CIFAR10, and vice versa for Flower. The third row show that enabling DP only adds $9\%$ extra simulation wall-clock time.
Figure 2: Speedup from scaling up number of processes per GPU in distributed simulations, while keeping the hardware resources fixed. As long as number of GPUs $\ll$ cohort size (unlike the blue line, where we use an unnecessarily large amount of GPUs given the size of model, users and cohort), the wall-clock time is monotonically decreasing when increasing number of models to train in parallel on the same GPU.
The number of processes per GPU, $p$, was tuned to minimize wall-clock time, and the results show that pfl-research needs many fewer training processes per GPU to saturate GPU utilization. Tables 1 and 2 demonstrate that before the release of pfl-research, the comparison of framework speed greatly depended on the choice of benchmarks.
The superior performance of pfl-research can be attributed to several design choices, some of which are unique to it or not commonly employed by other FL simulators, described previously in Section 3. Appendix D.4.2 show system metrics from training the CIFAR10 benchmark, giving more insight into how pfl-research is faster.
@@ 4.2 Scaling distributed training
Figure 3: Speedup from scaling up distributed simulations. Left panel: Sweep number of processes per GPU on CIFAR10, StackOverflow and FLAIR benchmarks and keep the hardware resources pinned. Right panel: Sweep the number of GPUs to train the StackOverflow benchmark, repeat for 1, 3 and 5 processes per GPU. Blue lines (tracked on left y-axis) show wall-clock time. Green lines (tracked in right y-axis) show total GPU hours for the same experiments. Note that the runs with >8 GPUs use multiple hosts and thus become slightly less efficient.
Figure 2 shows the wall-clock time when scaling up number of processes in distributed simulations. In pfl-research, the number of worker processes is decoupled from the number of GPUs available, allowing processes to share a GPU. Using more than one process per GPU, $p>1$, always has a positive effect, but the optimal number depends on the use-case (model size, cohort size). For CIFAR10 and StackOverflow [80], we can set $p=5$ and the wall-clock time is less than $0.43$ of that with $p=1$ with same resources, while for FLAIR we can only fit a maximum of $p=3$ on the GPU and the relative speedup is smaller.
The left panel of Figure 3 shows the wall-clock time when scaling up the number of GPUs on the StackOverflow benchmark with a cohort size of 400. The wall-clock time is reduced to 3 hours with just one GPU when scaling up number of worker processes to five, and the wall-clock time is reduced to 15 minutes when scaling up to 32 GPUs. Thus, pfl-research scales well with number of GPUs. The rate of improvement is expected to be reduced as we increase the total number of GPUs because there are fewer users per process to load balance. This is revealed by the total number of GPU hours to finish (green lines). To demonstrate a more compute intensive scenario that can be scaled up further, we increase cohort size to 50,000 and show results from scaling up in the right panel of Figure 3. We see that distributed simulations scale even better in this case, for example the increase in GPU hours when moving from 16GPUs to 32GPUs is only 3.6%.
@@ 4.3 Benchmarks for research
Table 3: Performance of FL algorithms on the pfl-research benchmark suite without DP. C10 stands for accuracy on CIFAR10, SO stands for perplexity on StackOverflow, FLR stands for mean average precision (mAP) on FLAIR. The last three columns are the perplexity scores for three LLM benchmark datasets: Stanford Alpaca (SA), Aya, and OpenAssistant (OA). If not mentioned, the natural or artificial non-IID partitioning is used for the dataset, and -IID means partitioning with fixed number of samples per client drawn IID. AdaFedProx is FedProx with adaptive $\mu$ from Appendix C.3.3 in [52]. Errors for LLM experiments are detailed in Table 12
@@
FedAvg [60] $70.37_{\pm.21}$ $69.80_{\pm.35}$ $60.87_{\pm.28}$ $65.08_{\pm.60}$ $61.83_{\pm.10}$ $3.61$ $4.02$ $6.39$
FedProx [52] $70.69_{\pm.33}$ $69.79_{\pm.35}$ $60.88_{\pm.28}$ $64.45_{\pm 1.13}$ $62.02_{\pm.11}$ $3.61$ $4.02$ $6.39$
AdaFedProx [52] $70.39_{\pm.33}$ $69.61_{\pm.45}$ $61.01_{\pm.34}$ $64.53_{\pm 1.01}$ $61.99_{\pm.11}$ $3.61$ $4.02$ $6.40$
SCAFFOLD [42] - $60.17_{\pm 1.29}$ $62.43_{\pm.09}$ $60.82_{\pm.12}$ $3.73$ $4.77$ $6.90$
Table 4: Performance of FL algorithms on the pfl-research benchmark suite with central DP. The same rules and notation as in Table 3 are used. G stands for Gaussian mechanism with accounting using privacy loss distribution [62, 22] and BMF stands for banded matrix factorization mechanism [20]. Errors for LLM experiments are detailed in Table 13.
@@
FedAvg G $68.77_{\pm.38}$ $66.70_{\pm.09}$ $69.11_{\pm.35}$ $65.24_{\pm.22}$ $61.49_{\pm.11}$ $3.62$ $4.19$ $6.52$
FedAvg BMF $69.21_{\pm.51}$ $67.07_{\pm.14}$ $62.47_{\pm.22}$ $65.55_{\pm.05}$ $61.64_{\pm.10}$ $3.62$ $4.20$ $6.53$
FedProx G $68.72_{\pm.34}$ $66.75_{\pm.21}$ $69.12_{\pm.34}$ $65.24_{\pm.17}$ $61.55_{\pm.11}$ $3.62$ $4.20$ $6.52$
AdaFedProx G $68.78_{\pm.48}$ $66.80_{\pm.09}$ $69.14_{\pm.34}$ $65.18_{\pm.16}$ $61.58_{\pm.08}$ $3.62$ $4.20$ $6.52$
SCAFFOLD G - $61.31_{\pm.24}$ $77.71_{\pm.10}$ $51.64_{\pm.44}$ $3.79$ $4.76$ $7.00$
The initial set of algorithms we benchmark is shown in Table 3 without DP and in Table 4 with central DP. All experiments are run for five times with different random seed and metrics are averaged. Evaluation is done on the validation data partitions of the original datasets without any federated splits. The common hyperparameters, such as cohort size and number of local epochs, were optimized using the FedAvg experiments. Meanwhile, specific hyperparameters for each algorithm were individually tuned. As a result, direct comparisons between the performance of different algorithms and FedAvg may not be conclusive. However, some observations based on the results include:
  • Banded matrix factorization mechanism, also known as DP-FTRL when applied to FL, significantly outperforms Gaussian mechanism with PLD moments accountant on StackOverflow, with 10% relative improvement.
  • SCAFFOLD, a popular algorithm according to citations, does not perform better than FedAvg on any of our benchmarks. [63] also report that SCAFFOLD consistently perform worse than FedAvg.
  • FedProx is known for performing well on heterogeneous data partitions, but this is only marginally reflected in the FLAIR benchmark results. It is worse than FedAvg baseline on FLAIR IID, but better on FLAIR with natural heterogeneous client partitions.
The breadth of algorithms is limited at this moment, but we believe we have provided robust FedAvg baselines and a flexible foundation for future work to expand the results in terms of algorithms, complementary datasets, better hyper-parameter tuning of algorithms for more fair comparison, and other evaluation metrics, e.g. amount of communicated bits.
LLM benchmarks
We consider three datasets for benchmarking LLM fine-tuning with pfl-research to demonstrate the scalability to larger models: Stanford Alpaca (SA), Aya and OpenAssistant (OA). Compared to existing FL LLM benchmarks [90, 49] where centralized datasets are artificially partitioned into users, Aya and OA have a natural user partition to simulate heterogeneity that reflects a more realistic production setting. SA contains 52,002 instruction-following data which is then artificially partitioned into users in an IID fashion. Aya contains a total of 204,112 human-annotated prompt-completion pairs, where each label has the identifier of the annotator. OA contains more than 120,000 conversational messages between human and AI assistant pairs. We collect 85,318 pairs of user inputs and assistant response with associated user identifier. We partition Aya and OA into users in a non-IID fashion using the provided user identifiers. Details of dataset partition and model fine-tuning are described in Appendix 11.
@@ 5 Future work
pfl-research currently provides a small collection of federated algorithms. Our primary goal has been to release a fast, modular and easy-to-use simulator which is useful for advancing research in this area. We will continuously benchmark new algorithms and methods, and will rely heavily on the community for contributing implementations of new research.
We plan to carefully curate more benchmarks while maintaining a diverse set of domains for the benchmarks. All benchmark results reported in Section 4.3 used PyTorch, and releasing similar benchmarks in TensorFlow is work in progress. We recognize that all our current benchmarks use cross-device FL. Cross-silo FL can be simulated with pfl-research, but we do not provide benchmarks for this scenario yet.  [^2] We refer to https://github.com/apple/pfl-research/blob/main/pfl/data/sampling.py#L143 for more information about cross-silo simulation in pfl-research.
The aggregator interface can be extended to make a client-server architecture with a dedicated aggregator and coordinator process, which might be required for testing certain research ideas not limited to synchronous FL. We also do not provide interfaces for implementing PFL with real edge devices. This is outside the scope of pfl-research at this time.
In some scenarios, such as large-scale cross-silo PFL simulations or training foundation models, it may be useful to train a local model for a single user (silo) on multiple GPUs. Implementation of model parallelism in pfl-research remains for future work.
Some benchmarking related to system performance left for future work are on the concept of weak scaling—to investigate the efficiency of the distributed setup as the workload and resources scale proportionally. Figure 3 had a fixed cohort size and didn’t increase the cohort sizes proportionally.
@@ Appendix A Preliminaries
Appendix B Detailed system design
B.1 API Design
pfl-research is object-oriented Python 3 code, using the following class hierarchies.
Dataset    FederatedDataset parameterizes how to partition, load, and pre-process the data of an individual user, and iterating over one produces per-user Datasets. The interfaces are flexible enough to support loading data from any storage using any preprocessing library. pfl-research also provides integration with tf.data.Dataset and torch.utils.data.Dataset (from here on referred to as "native datasets"). If the datasets of clients are small enough such that several can be loaded into memory simultaneously, a native dataset can be defined on the same level as FederatedDataset, returning the full data tensors of one user when requesting the next tensors. This is applicable in cross-device FL, where user datasets are typically small enough to fit into memory. If client datasets are large and loading the whole dataset at once is inefficient or requires too much memory, e.g. in large-scale cross-silo simulations, one native dataset can represent one client, with asynchronously loaded batches.
Model    A Model is an adapter for connecting a TensorFlow or PyTorch models to the simulator. Additionally, the Model class can be extended to implement non-neural-network models. pfl-research provides an implementation of federated gradient boosted decision trees (GBDTs) [27] and federated Gaussian mixture models (GMMs).
Algorithm    The interface for implementing federated algorithms is FederatedAlgorithm. We generalize the responsibilities of an algorithm to be:
Aggregator    An Aggregator is a lightweight interface for (1) accumulating model updates from clients and (2) reducing partial aggregated model updates across worker processes. The default is accumulating with a sum, and reducing across workers with an all-reduce sum; see Section 3.1 for further explanation.
Backend    A Backend consumes the federated datasets, postprocessors (see below), and aggregator to implement the simulation control flow. Since pfl-research was released for running simulations, only the concrete backend currently included is SimulatedBackend.
Postprocessor    A Backend can invoke a series of Postprocessors. These can implement features such as compression, sparsification, and DP mechanisms by modifying local model updates as a postprocessing step, and optionally by postprocessing the aggregated model update before it is sent back to the algorithm for the central model update procedure. The server-side postprocessing steps are run in reversed order. It is essential to be mindful of the order of postprocessors, e.g. DP mechanisms commonly fit best as the last postprocessor for local updates so that no further modification of the sensitivity of the model update happens after clipping.
Callback    TrainingProcessCallback provides a hook into the central training loop. A callback does not have access to the data or model update and is called only after the central model has been updated, hence it should not alter the learning in any way. Some callback implementations that pfl-research provides are: fault-tolerant training procedure, evaluation on central dataset, exponential moving average of model, stopping criterion, reporting intermediate results (csv files, TensorBoard and Weights & Biases [^3] https://wandb.ai/), and profiling tools.
Hyperparameters    There are many hyperparameters to tune in (P)FL experiments, and there has been research on constructing schedules and adaptive approaches to simplify the hyperparameter tuning process [5, 70, 47]. In pfl-research, many hyperparameters of local model training or the algorithm can be expressed as either simple types (constant for the experiment) or as a HyperParam (variable across iterations). At the start of each central iteration, the algorithm requests the current value of any HyperParam instances to set as the static value for that iteration only. A concrete hyperparameter class can in turn extend Postprocessor or TrainingProcessCallback to hook into the training loop and adapt its value based on arbitrary rules.
Metrics    Metrics have conflicting implementations in various FL frameworks. We define two types below. See a concrete example for each kind in Appendix B.4.
  • Per-user metrics — Each client produces the metric, and the aggregation over clients’ results is the average of such metrics. This is more relevant when the performance of each client is of interest, e.g. in personalization and meta-learning.
B.2 FL simulation
Algorithm 1 Generalized federated learning simulation.
@@ B.2 FL simulation
2       /* The algorithm generates a set of contexts $\mathcal{C}$ which are instructions on how to gather results from cohorts. May modify model in preparation for this iteration. */ ($\mathcal{C},\theta_{t}^{\prime})\leftarrow$ $alg$.get_next_central_contexts($\theta_{t}$, $t$)  if $\mathcal{C}=\emptyset$ then
3             // Algorithm signaled that training should end. break
4      $\mathcal{S}\leftarrow\emptyset$ /* Sample a cohort and get aggregated results for each context. This is done by the Backend in the implementation. */ foreach central context $c_{i}\in\mathcal{C}$ do
5             $D\leftarrow$ $\mathcal{F}$, $c_{i}$.population)  foreach worker $w$ in parallel do
6                   $\Delta^{w}_{t,i}\leftarrow null$foreach user dataset $d_{u}\in\mathrm{distributed\_sample}(D,w,c_{i}.\mathrm{cohort\_size)}$ do
7                         $(\Delta_{t,i}^{u},aux_{u})\leftarrow alg$.simulate_one_user($\theta_{t}^{\prime}$, $d_{u}$, $c_{i}$)  if $\Delta_{t,i}^{u}\neq null$ then
8                               // The results from local optimization is passed through all postprocessors, which may manipulate the statistics and use auxiliary non-private information from the user. foreach postprocessor $p\in\mathcal{P}$ do
9                                     $\Delta_{t,i}^{u}\leftarrow p$.postprocess_one_user($\Delta_{t,i}^{u}$, $aux_{u}$)
10                              $\Delta^{w}_{t,i}\leftarrow agg$.accumulate($\Delta^{w}_{t,i}$, $\Delta_{t,i}^{u}$)
11
12                   // Synchronize aggregates from all workers. $\Delta_{t,i}\leftarrow agg$.worker_reduce($\Delta^{w}_{t,i}$, $c_{i}$, $w$)
13            foreach postprocessor $p\in\mathrm{reversed}(\mathcal{P})$ do
14                   $\Delta_{t,i}\leftarrow p$.postprocess_server($\Delta_{t,i}$, $c_{i}$)
15            $\mathcal{S}\leftarrow\mathcal{S}\cup\{\Delta_{t,i}\}$
16       /* The algorithm consumes the set of aggregated statistics from all queried cohorts and produces a new central model. */ $\theta_{t+1}\leftarrow alg$.process_aggregated_statistics_all_contexts($\mathcal{C}$, $\mathcal{S}$, $\theta^{\prime}_{t}$)  foreach callback $b\in\mathcal{B}$ do
17             // A callback may force stop training, e.g. early stopping. $stop$ |= $b$.after_central_iteration($\theta_{t+1}$, $t$)
18      $t\leftarrow t+1$
19until $stop$
An algorithm runs for a number of central iterations, which is either determined by the algorithm or interrupted by a callback. A context $c_{i}\in\mathcal{C}$ is a recipe for how to run the next central iteration. It contains parameters for local model training and evaluation, the state of the algorithm, dataset target criteria, and parameters for other steps of the algorithm. A federated algorithm $alg$ operates on a group of federated datasets $\mathcal{F}$ and each context $c_{i}$ targets one federated dataset $D\in\mathcal{F}$.